OpenStudy (anonymous):

Find the value of K such that the equation X^2+Y^2+6x-10y+K=0 represents a circle with a radius of 5.

OpenStudy (amistre64):

(x^2 +6x) + (y^2 -10y) +K= 0 Complete the squares of your x and y: (x^2 +6x +a) + (y^2 -10y +b) = 0; K= a+b (x^2 +6x +(6/2)^2) + (y^2 -10y +(10/2)^2) = 0 K = 3 + 5 = 8 If I remember it correctly :)

OpenStudy (anonymous):

Thank you!

OpenStudy (amistre64):

I would need to double check that to be sure if I am right..

OpenStudy (amistre64):

I was close. K needs to equal (3)^2+(5)^2 = 9+25 = 34. And the radius of 5 would equal (5)^2 = 25. K=34-25 = 11 is the correct answer.

OpenStudy (anonymous):

Thank you very much!

OpenStudy (amistre64):

Dont thank me yet, I forgot how to subtract :) 34-25 = 9. Lets go with that :)

OpenStudy (anonymous):

I hope I can thank you now. THANK YOU!

OpenStudy (amistre64):

Yep, that one works. Youre Welcome :)